what is the equation of the line that contains the points with(x,y)coordinates(-3,7)and(5,-1)

Answer by checkley71(8403)

slope=(-1-7)/(5-3)=-8/2=-4
7=-4*3+b
7=-12+b
b=7+12
b=19
y=-4x+19